OpenStudy (anonymous):

Topic:the fundamental principle of counting How many ways can 3 cars be parked in 4 parking spaces?

OpenStudy (anonymous):

does the order matter?

OpenStudy (anonymous):

yah i think your correct 4c3? ryt? sorry i just chatted on mathematic i didnt see you reply sorry

OpenStudy (anonymous):

ops 4p3? or 4c3?

OpenStudy (raden):

4P3 = 4!/(4-3)! = 4!/1! = 4.3.2.1/1 = 24 ways

OpenStudy (anonymous):

thanks eddie and raden

OpenStudy (anonymous):

i think we can assume these are 3 distinct cars. there is 4 spots that the first car can go in, then 3 remaining spots for the second car, then 2 remaining spots to choose from for the third car. by fundamental counting principle we get 4x3x2

OpenStudy (anonymous):

another 1 i got my assignment here i post another i closed this one

OpenStudy (raden):

thought one

OpenStudy (anonymous):

do you agree?